OpenStudy (my4babiesandme):

solve (r-8s+3)+(7r+7s)+(s+2)

myininaya (myininaya):

r+7r-8s+7s+s+3+2 (1+7)r+(-8+7+1)s+(3+2)

myininaya (myininaya):

1+7=8 -8+7+1=0 3+2=5

OpenStudy (my4babiesandme):

its supposed to have r and s in the answer

myininaya (myininaya):

yes it does replace 1+7 with 8 replace -8+7+1 with 0 replace 3+2 with 5 in the expression i have above

myininaya (myininaya):

it is

myininaya (myininaya):

r+7r-8s+7s+s+3+2 (1+7)r+(-8+7+1)s+(3+2) 1+7 is 8 -8+7+1 is 0 3+2 is 5 8r+0s+5 8r+5

myininaya (myininaya):

what part are you getting stuck on

OpenStudy (my4babiesandme):

i got it thanks